# Random Turán Problems for Hypergraph Expansions

### v1

@@ 1.2 Main Results: Balanced Supersaturation
This definition together with a standard application of hypergraph containers yields the following result. We note that particular cases of this result have appeared throughout earlier works on this subject, and we expect that this general version of the result will be quite convenient to use in future study of random Turán problems (see also 4.1 for general lower bounds).
@@
We emphasize that condition (b) here is quite mild: for any $\tau$ such that (a) holds, we can always apply 1.5 after replacing $\tau$ with the new function $\tau^{\prime}(n,m):=1+\max_{m^{\prime}\geq m}\tau(n,m^{\prime})$ since (a) continues to hold if we increase $\tau$.
With 1.5 in mind, our goal now will be to prove a set of Lifting Theorems which show that expansions $F^{(r)}$ are balanced whenever $F$ is. Our first result of this form can be seen as a generalization of the proof of [13, Theorem 5.1, Part 1]. While the statement is rather technical, we emphasize that it essentially just says that if $F$ is $(M,\gamma,\tau)$-balanced for some reasonable $M,\gamma,\tau$, then every expansion $F^{(r)}$ is also balanced in some appropriate way.
@@
Our second Lifting Theorem, which can be viewed as an extension of the proof of Theorem 2.6 and Theorem 2.7 in [15], gives stronger bounds than 1.6 when going from $F^{(r-1)}$ to $F^{(r)}$ but can only be applied when $r$ is sufficiently large. To state this, we require a few extra definitions.
@@
Given an $r$-graph $F$ with at least 2 edges, we denote its $r$-density by
$d_{r}(F):=\max_{F^{\prime}\subseteq F,\ |F^{\prime}|\geq 2}\left\{\frac{|F^{\prime}|-1}{v(F^{\prime})-r}\right\}.$
@@
1.7 turns out to be essentially at least as strong as 1.6 whenever it applies; see the arXiv only appendix for more on this. On the other hand, 1.7 is incapable of proving bounds on $F$ such as $K_{s,t}^{(r)}$ with $r\leq s$ since the tight tree condition does not hold here . As such, our forthcoming work [16] on improving the bound on $r$ in 1.3 will require us to work exclusively with 1.6.
By using 1.7, we will show that if $F$ is an $r_{0}$-graph which has “optimal” balanced supersaturation (see 3.1) and Turán number $O(n^{r_{0}-1})$ (i.e. one whose extremal construction has the same size as that of an $r_{0}$-uniform star), then every expansion of $F$ has essentially optimal bounds for the random Turán problem. For minor technical reasons we restrict our result to $F$ with maximum degree $\Delta(F)$ not too large or small.
@@
Organization. The rest of our paper is organized as follows. We prove Theorems 1.6 and 1.7 in Section 2, which is the bulk of the technical work of this paper. We then prove several results related to optimal balanced supersaturation in Section 3 and general random Turán results in Section 4, culminating in a proof of 1.8. Finally, we apply our Lifting Theorems to give short proofs of Theorems 1.2 and 1.4 in Section 5, with the proof of 1.3 being deferred to an arXiv only appendix.
@@ 2.1 Proof Sketch
In this section we prove our two general results Theorems 1.6 and 1.7 about lifting balanced supersturation from $F$ to the expansions $F^{(r)}$. To illustrate our methods, we informally consider the toy problem of finding copies of $C_{4}^{(3)}$ in a 3-graph $H$ where every pair of vertices in $H$ has degree either 0 or $d$. In this setting we use one of two approaches to build our copies of $C_{4}^{(3)}$.
@@
We informally refer to the first approach described here as “Shadow Expansion” and the second as “Greedy Expansion”. We will prove 1.6 by using only shadow expansion, and we will prove 1.7 by combining both of these methods.
2.2 Shadow Expansion: Proof of 1.6
@@
We now have all the tools we need to prove our first main result 1.6, which we restate below for convenience.
@@
Let $H_{[r_{0}]}$ be the $r_{0}$-graph on $\bigcup_{i\in[r_{0}]}V_{i}$ whose edges are all $r_{0}$-shadows of edges of $H^{\prime\prime}$ in $\bigcup_{i\in[r_{0}]}V_{i}$. Let $\hat{n}=|\bigcup_{i\in[r_{0}]}V_{i}|$ and let $\hat{m}=|H_{[r_{0}]}|$. Our aim now will be to apply 2.1 to $H_{[r_{0}]}$, and to this end we need to effectively lower bound $\hat{m}$ and, in view of (10), we will also need to show that $d:=\Delta_{[r_{0}]}(2r\log n)^{-2^{r}}$ is at least $2v(F^{(r)})n^{r-r_{0}-1}$. Towards both of these goals, we observe by (10) that
@@
2.3 Greedy Expansion: Proof of 1.7
We now prove our second result 1.7 which lifts balancedness from $(r-1)$-graphs $F$ to $F^{(r)}$ whenever $F^{(r)}$ is contained in a tight $r$-tree. We begin by introducing two lemmas that will be useful. The following says that every $r$-partite $r$-graph contains a large subgraph such that either all $(r-1)$-shadows have large codegrees or all $(r-1)$-shadows in some $(r-1)$ parts have small (but not too small) and regular codegrees.
@@
The following lemma comes from [14, Lemma 2.2] and is proven in essentially the same way as described in the second approach of Section 2.1.
@@
We are now ready to prove 1.7, which we restate below for convenience.
@@ 3 Optimal Balanced Supersaturation
Our main goal of this section is to prove a result (3.4) giving a sufficient condition for expansions to have optimal balanced supersaturation, which will be the main ingredient for proving 1.8. We will need two lemmas regarding expansions, the first of which shows that $r$-densities play nicely with expansions.
@@
$\frac{1}{d_{r_{0}}(F)}\leq\frac{v(F^{\prime\prime})-r_{0}}{|F^{\prime\prime}|-1}\leq\frac{v(F^{\prime})-r_{0}-(r-r_{0})|F^{\prime}|}{|F^{\prime}|-1}=\frac{1}{d_{r}(F^{(r)}}-r+r_{0},$
@@
It is straightforward to check that properties (a), (b), and (c) of 1.7 are satisfied, and we have that $F^{(r)}$ is a spanning subgraph of a tight tree due to 3.3 applied with $k=r_{0}$. Let $C$ be the constant guaranteed to exist from 1.7. Let $M_{r}(n)=Cn^{r-1}$, let $\gamma_{r}(n)=\gamma_{r-1}(n)\cdot C\log n=C(\log n)^{C_{r-1}+1}$, and let
@@
We claim that property (d) of 1.7 holds. Indeed, the lower bound $A(n,m)\geq mn^{1-r}$ holds trivially since this is one of the terms in the maximum. The upper bound $A(n,m)\leq\frac{m}{M_{r-1}(n)C\log n}=\frac{m}{C_{r-1}n^{r-2}\cdot C\log n}$ holds for the second term of $A$ provided (${n\choose r-1}\geq C_{r-1}n^{r-2}$, which holds for $n$ sufficiently large, and it holds for the first term provided
@@
Because $d_{r}(F^{(r)})<1$ by 3.2, this inequality is hardest to satisfy at the smallest possible value of $m\geq Cn^{r-1}$, and one can verify this holds for $n$ sufficiently large through the inequality $-d_{r}(F^{(r)})+2-r<2-r$.
@@
where this last step used $\tau_{r-1}=n^{r-1-\frac{1}{d_{r-1}(F)}}(\log n)^{C_{r-1}}=n^{r-\frac{1}{d_{r}(F^{(r)}}}(\log n)^{C_{r-1}}$ by 3.2. We conclude by 1.7 that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced. ∎
@@ 4 General Random Turán Results
Here we prove a set of general lower and upper bounds for the random Turán problem, special cases of which have appeared throughout the literature in previous papers. We begin with some simple bounds.
Lemma 4.1.
@@
We now move onto our upper bound for the random Turán problem 1.5. Our proof uses the hypergaph container method, developed independently by Balogh–Morris–Samotij [1] and Saxton–Thomason [20], which is fundamental in the study of random Turán problems. Here we make use of the following simplified version of Theorem 1.5 in [9]:
@@
We now proving 1.5, which we restate below for convenience.
@@
Let $n$ be sufficiently large and $m\geq M(n)$. Note that by 3.1 we have $m\geq\mathrm{ex}(n,F)=\Omega(n)$ since $r,|F|\geq 2$, and in particular for $n$ sufficiently large we have
@@
The upper bound comes from applying 1.5 with $m=M(n)$, and the lower bound comes from 4.1. ∎
We now prove 1.8 which we recall below.
@@
By 3.4 (whose hypothesis is exactly the same as that of 1.8), we find for all $r\geq r_{0}$ that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced with $M(n)=\Theta(n^{r-1})$, $\gamma=(\log n)^{\Theta(1)}$ and $\tau=n^{r-\frac{1}{d_{r}(F)}}(\log n)^{\Theta(1)}$.
@@
We claim that $\mathrm{ex}(n,F^{(r)})=\Theta(n^{r-1})$. Indeed, the lower bound comes from a star (i.e. every edge containing a given vertex $v\in[n]$), which is $F$-free since $\Delta(F)<|F|$ (meaning no vertex of $F^{(r)}$ is contained in every edge of $F^{(r)}$); and the upper bound comes from $M(n)=O(n^{r-1})$ and 3.4. We conclude the result by 4.4. ∎
5 Applying our Results

### v2

@@ 2.1 General Tools for Random Turán Problems
This definition together with a standard application of hypergraph containers yields the following result.
@@
We emphasize that condition (b) here is quite mild: for any $\tau$ such that (a) holds, we can always apply 2.6 after replacing $\tau$ with the new function $\tau^{\prime}(n,m):=1+\max_{m^{\prime}\geq m}\tau(n,m^{\prime})$ since (a) continues to hold if we increase $\tau$.
With 2.6 in hand, we see that to prove our main results we need only show that Sidorenko expansions $F^{(r)}$ are balanced in some appropriate sense. To do this, we establish several “Lifting Theorems” which show that expansions $F^{(r)}$ are balanced whenever the base hypergraph $F$ is. These set of Lifting Theorems provide significant generalizations to similar results of Mubayi and Yepremyan [25] and Nie [27] and can be viewed as the main technical innovation of the present work.
We begin with the first result in this direction. While the statement is rather technical, we emphasize that it is simply saying that if $F$ is $(M,\gamma,\tau)$-balanced for some reasonable choice of $M,\gamma,\tau$, then every expansion $F^{(r)}$ is also balanced in some appropriate way.
@@
Our second Lifting Theorem gives stronger bounds than 2.7 when going from $F^{(r-1)}$ to $F^{(r)}$ but can only be applied when $r$ is sufficiently large. To state this, we require an additional definition.
Definition 4.
@@
2.8 turns out to be essentially at least as strong as 2.7 whenever it applies; see the arXiv-only appendix of this paper for more on this.
By using 2.8, we will be able to show that if $F$ is an $r_{0}$-graph which has “optimal” balanced supersaturation (see 4.1) and Turán number $O(n^{r_{0}-1})$ (i.e. one whose extremal construction has the same size as that of an $r_{0}$-uniform star), then every expansion of $F$ has essentially optimal bounds for the random Turán problem in the following sense.
@@
Organization. The rest of our paper is organized as follows. We prove Theorems 2.7 and 2.8 in Section 3, which is the bulk of the technical work of this paper. We then prove several results related to optimal balanced supersaturation in Section 4 and general random Turán results in Section 5, culminating in a proof of 2.9. Finally, we apply our Lifting Theorems to prove our main results on $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ for various Sidorenko $F$ in Section 6.
@@ 3 Balanced Supersaturation for Expansions
In this section we prove our two general results Theorems 2.7 and 2.8 about lifting balanced supersaturation from $F$ to the expansions $F^{(r)}$. For this, given an $r$-graph $H$ we say that a set of vertices $S$ with $|S|=r_{0}$ is an $r_{0}$-shadow of $H$ if there exists an edge of $H$ containing the set $S$.
@@ 3.1 Proof Sketch
To illustrate our methods, we informally consider the toy problem of finding copies of $C_{4}^{(3)}$ in a 3-graph $H$ where every pair of vertices in $H$ has degree either 0 or $d$. In this setting we use one of two approaches to build our copies of $C_{4}^{(3)}$.
@@
We informally refer to the first approach described here as “Shadow Expansion” and the second as “Greedy Expansion”. We will prove 2.7 by using only shadow expansion, and we will prove 2.8 by combining both of these methods.
3.2 Shadow Expansion: Proof of 2.7
@@
We now have all the tools we need to prove our first main result 2.7, which we restate below for convenience.
@@
Let $H_{[r_{0}]}$ be the $r_{0}$-graph on $\bigcup_{i\in[r_{0}]}V_{i}$ whose edges are all $r_{0}$-shadows of edges of $H^{\prime\prime}$ in $\bigcup_{i\in[r_{0}]}V_{i}$. Let $\hat{n}=|\bigcup_{i\in[r_{0}]}V_{i}|$ and let $\hat{m}=|H_{[r_{0}]}|$. Our aim now will be to apply 3.1 to $H_{[r_{0}]}$, and to this end we need to effectively lower bound $\hat{m}$ and, in view of (10), we will also need to show that $d:=\Delta_{[r_{0}]}(2r\log n)^{-2^{r}}$ is at least $2v(F^{(r)})n^{r-r_{0}-1}$. Towards both of these goals, we observe by (10) that
@@
3.3 Greedy Expansion: Proof of 2.8
We now prove our second result 2.8 which lifts balancedness from $(r-1)$-graphs $F$ to $F^{(r)}$ whenever $F^{(r)}$ is contained in a tight $r$-tree. We begin by introducing two lemmas that will be useful. The following says that every $r$-partite $r$-graph contains a large subgraph such that either all $(r-1)$-shadows have large codegrees or all $(r-1)$-shadows in some $(r-1)$ parts have small (but not too small) and regular codegrees.
@@
The following lemma comes from [26, Lemma 2.2] and is proven in essentially the same way as described in the second approach of Section 3.1.
@@
We are now ready to prove 2.8, which we restate below for convenience.
@@ 4 Optimal Balanced Supersaturation
Our main goal of this section is to prove a result (4.4) giving a sufficient condition for expansions to have optimal balanced supersaturation, which will be the main ingredient for proving 2.9. We will need two lemmas regarding expansions, the first of which shows that $r$-densities play nicely with expansions.
@@
$\frac{1}{d_{r_{0}}(F)}\leq\frac{v(F^{\prime\prime})-r_{0}}{|F^{\prime\prime}|-1}\leq\frac{v(F^{\prime})-r_{0}-(r-r_{0})|F^{\prime}|}{|F^{\prime}|-1}=\frac{1}{d_{r}(F^{(r)})}-r+r_{0},$
@@
It is straightforward to check that properties (a), (b), and (c) of 2.8 are satisfied, and we have that $F^{(r)}$ is a spanning subgraph of a tight tree due to 4.3 applied with $k=r_{0}$. Let $C$ be the constant guaranteed to exist from 2.8. Let $M_{r}(n)=Cn^{r-1}$, let $\gamma_{r}(n)=\gamma_{r-1}(n)\cdot C\log n=C(\log n)^{C_{r-1}+1}$, and let
@@
We claim that property (d) of 2.8 holds. Indeed, the lower bound $A(n,m)\geq mn^{1-r}$ holds trivially since this is one of the terms in the maximum. The upper bound $A(n,m)\leq\frac{m}{M_{r-1}(n)C\log n}=\frac{m}{C_{r-1}n^{r-2}\cdot C\log n}$ holds for the second term of $A$ provided (${n\choose r-1}\geq C_{r-1}n^{r-2}$, which holds for $n$ sufficiently large, and it holds for the first term provided
@@
Because $d_{r}(F^{(r)})<1$ by 4.2, this inequality is hardest to satisfy at the smallest possible value of $m\geq Cn^{r-1}$, and one can verify this holds for $n$ sufficiently large through the inequality $-d_{r}(F^{(r)})+2-r<2-r$.
@@
where this last step used $\tau_{r-1}=n^{r-1-\frac{1}{d_{r-1}(F)}}(\log n)^{C_{r-1}}=n^{r-\frac{1}{d_{r}(F^{(r)}}}(\log n)^{C_{r-1}}$ by 4.2. We conclude by 2.8 that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced. ∎
@@ 5 General Random Turán Results
Here we prove a set of general lower and upper bounds for the random Turán problem, special cases of which have appeared throughout the literature in previous papers. We begin with proving the simple bounds of 2.5, which we restate here for the reader’s convenience.
Proposition 2.5.
@@
We now move onto our upper bound for the random Turán problem 2.6. Our proof uses the hypergaph container method, developed independently by Balogh–Morris–Samotij [1] and Saxton–Thomason [32], which is fundamental in the study of random Turán problems. Here we make use of the following simplified version of Theorem 1.5 in [21]:
@@
We now proving 2.6, which we restate below for convenience.
@@
Let $n$ be sufficiently large and $m\geq M(n)$. Note that by 4.1 we have $m\geq\mathrm{ex}(n,F)=\Omega(n)$ since $r,|F|\geq 2$, and in particular for $n$ sufficiently large we have
@@
The upper bound comes from applying 2.6 with $m=M(n)$, and the lower bound comes from 2.5. ∎
We now prove 2.9 which we recall below.
@@
By 4.4 (whose hypothesis is exactly the same as that of 2.9), we find for all $r\geq r_{0}$ that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced with $M(n)=\Theta(n^{r-1})$, $\gamma=(\log n)^{\Theta(1)}$ and $\tau=n^{r-\frac{1}{d_{r}(F)}}(\log n)^{\Theta(1)}$.
@@
We claim that $\mathrm{ex}(n,F^{(r)})=\Theta(n^{r-1})$. Indeed, the lower bound comes from a star (i.e. every edge containing a given vertex $v\in[n]$), which is $F$-free since $\Delta(F)<|F|$ (meaning no vertex of $F^{(r)}$ is contained in every edge of $F^{(r)}$); and the upper bound comes from $M(n)=O(n^{r-1})$ and 4.4. We conclude the result by 5.4. ∎
6 Proofs of Main Results